Q.Choose the correct answer.
Q. For what value of r all solutions of the equation
u^''-(2r-1)u^'+r(r-1)u=0 tend to zero as x→∞?
a)r=1
b)r<0
c)r>0
d)r<1
1
Expert's answer
2019-07-09T12:19:25-0400

The general solution of given differential equation is "u(x)=C_1\cdot exp((r-1)x)+C_2\cdot exp(rx)". Hence "u(x)" tends to zero as "x \to \infty" when "r<0" . So the answer is b)"r<0."